# Quantum Otto heat engine and quantum Mpemba effect in quasiperiodic systems

### v1

@@ IV Dissipation driven delocalization-localization transition
Besides, some studies suggest this transition is independent of the system’s initial properties. Alternative viewpoints argue that the system’s initial attributes can influence the transition. Different initial property may either shorten or extend the threshold required for the delocalization-localization shift to occur. We aim to study whether the dependence of system’s initial property will also occur in this system with diagonal and off-diagonal quasiperiodic modulations and to study how different initial properties affect the delocalization-localization transition of the steady state as well.
@@
Upon introducing this bond dissipation Lindblad operator $L_{n}=c^{{\dagger}}_{n}c_{n+\ell}$, generated following the ones in Ref.Yusipov et al. (2017); Vershinina, O. S. et al. (2017); Yusipov et al. (2018); Vakulchyk et al. (2018); Jiang et al. (2024); Yang et al. (2024); Liu et al. (2024), then the time-evolution of the density matrix $\rho$ is dictated by the following Lindblad master equation
@@
Here, the dissipator $\mathcal{D}\left[L_{n}\right]$ is defined as $\mathcal{D}\left[L_{n}\right]=L_{n}\rho L^{{\dagger}}_{n}-\{L^{{\dagger}}_{n}L_{n},\rho\}/2$, and $\gamma_{n}=\gamma\cos({2\pi\alpha n})$ represents the tunably quasiperidoic bond dissipation strength with the incommensurate parameter $\alpha=\frac{\sqrt{5}-1}{2}$. The rationale for our selection of the quasiperiodic form for the key dissipation operator stems from the consideration of the following aspect: the initial key dissipation operator is made up of four terms, with two being dephasing terms Longhi (2024a) and the remaining two being hopping terms. As indicated in Ref. Liu et al. (2024), the dephasing effect fails to induce the delocalization-localization transition of steady states. It has been theoretically and experimentally confirmed that quasiperiodic modulation is capable of causing localization Roati et al. (2008). Moreover, this bond dissipation operator can be experimentally realized by introducing the auxiliary lattice Diehl et al. (2008); Kraus et al. (2008); Diehl et al. (2010, 2011); Bardyn et al. (2013). Owing to these factors, we utilized such a dissipative operator. This evolution equation can also be equivalently expressed as
@@
With $L=100$ and $\ell=1$, the delocalization phase diagrams for $V=0$, $V=0.5t$, $V=t$, and $V=2t$ are plotted in Figs. 5(a)- 5(d), respectively. The phase diagram is obtained by analyzing the position of the centroid in a steady state. We define a steady state where the centroid position is no more than five times the lattice constant from the system boundary as a localized state; otherwise, it is a delocalized state. The reason for choosing such a five-fold distance as the criterion is based on the experimental fact that in Roati et al. (2008), the authors reproduced a one-dimensional AA model in the cold atom optical lattice experimental system and observed the Anderson localization phase transition, where the wave packet broadening of the localized state was approximately $10$ times the lattice constant Roati et al. (2008). Therefore, it is feasible to distinguish the delocalization-localization transition of the steady state through the position of the center of mass.
@@
As can be seen from Figs. 5(a)-5(c), when $t_{1}$ is less than $0.5t$, there is a large parameter range that makes all the steady states delocalized. This indicates that under these parameters, the steady state can retain the initial delocalized property of the system. As $\gamma$ or $t_{1}$ increases, the steady state gradually changes from a delocalized state to a localized state, meaning that the properties of the steady state can be regulated by tuning $\gamma$ and $t_{1}$. Meanwhile, we observe that as $V$ gradually increases, the range of the delocalized region gradually shrinks. It is indicated that adjusting the parameter $V$ can also regulate the localization property of the steady state. For instance, when V is at a relatively small value, it can be seen from 5(a) and 5(b) that when $t_{1}>t$, there are still delocalized states within it, resulting in the reentrance delocalization phenomenon. Moreover, the value of $V$ is smaller, making the delocalization region larger. When $V$ is relatively large, it can be seen from Figs. 5(c) and 5(d) that the delocalized region disappears. In particular, we find that when $V=2t$, even if the initial system is in the localized phase, the steady state is still delocalized (as can be seen from Fig. 5(d), the delocalized state exists within a small range of $t_{1}$ and $\gamma$).
V quantum Mpemba effect and starting-line hypothesis
@@
Regarding to the delocalization-localization transition in open quasiperiodic system, the focus on this topic mainly lies in the localization properties of steady states. For an anomalous dynamic process: the relaxation rate of a system whose initial state is further from the equilibrium state is actually faster than that of a system whose initial state is closer to the equilibrium state, that is, the quantum Mpemba effect, remains to be studied and discussed.
@@
Figure 6: (Color Online) Trace distance $d_{\rm Tr}$ versus evolution time $\tau$. (a) $V=0$, $t_{1}=1.2t$, and $\gamma=0.5t$. (b) $V=0$, $t_{1}=1.7t$, and $\gamma=0.8t$. (c) $V=0.2t$, $t_{1}=1.2t$, and $\gamma=0.2t$. (d) $V=0.2t$, $t_{1}=1.5t$, and $\gamma=0.2t$. (e) $V=0.2t$, $t_{1}=0.8$, $\gamma=0.2t$. (f) $V=0.5t$, $t_{1}=0.5$, $\gamma=0.2t$. The blue regions denote the delocalization phases. The yellow regions denote the localization phases. Other parameter is $L=100$.
@@
We employ the density matrix $\rho(\tau)$ that describe the evolution of the system over time and the steady state density matrix $\rho_{ss}$ to calculate the standard distance between the matrices, this is, the trace distance, to characterize the quantum Mpemba effect. The trace distance $d_{\rm Tr}$ is defined as
@@
where $T=\rho(\tau)-\rho_{ss}$. Under Lindbladian dynamics, this quantity displays monotonic behavior over time. Notably, it has already been utilized as an experimental probe to investigate both the strong quantum Mpemba effect and its inverse counterpart Aharony Shapira et al. (2024); Zhang et al. (2025).
Prior to investigating the quantum Mpemba effect, we first introduce the ‘starting-line hypothesis’. This hypothesis posits the following: for an initial state whose energy is closest to the steady-state energy, if the distance between its center of mass and the steady-state center of mass is larger than that of other initial states with energies further from the steady-state energy, then the quantum Mpemba effect will arise. Before verifying this hypothesis, we can draw a simple inference regarding which initial states are capable of triggering the quantum Mpemba effect. For quantum states with extended properties, the position of their center of mass is consistently situated at the center of the system. In contrast, for quantum states with localized and critical properties, the positions of their centroids are not fixed at the center of system; instead, they can be distributed at other locations within the system. As a result, when quantum states with localized and critical properties are chosen as initial states, they will encounter the ‘starting-line discrepancy’ a condition that gives rise to the quantum Mpemba effect. On the other hand, for initial states with extended properties, their centroids all align at the same ‘starting-line’ (i.e., the center of system), meaning the quantum Mpemba effect will not occur.
@@
We take three different types of system eigenstates as the initial states and calculated the evolution of trace distance over time, as shown in Figures 6(a)-6(f). The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$. Specifically, the initial state corresponding to the red curve is the eigenstate of the system that is closest to the steady-state energy. Figures 6(a) and 6(b) show the results of the local state as the initial state. The centroid positions corresponding to the initial state (rounded off, the same below) are respectively: $n=\{60,49,39\}$ and $n=\{60,49,41\}$. Figures 6(c) and 6(d) show the results of delocalized states with critical properties as the initial states, where the corresponding centroid positions of the initial states are respectively: $n=\{63,61,30,83,49\}$ and $n=\{66,63,15,59,71\}$. Figures 6(e) and 6(f) present delocalized states with critical properties as the results of initial states, where the corresponding centroid positions of the initial states are respectively: all $n=50$ and all $n=51$. We can observe that for the initial states with extended properties, the dynamical evolution behavior does not exhibit the quantum Mpemba effect, while for the initial states with critical and localized properties, the dynamical evolution behavior will exhibit the quantum Mpemba effect. From the perspective of the centroid position of the initial state, when the Mpemba effect occurs, the wave packet of the initial state is relatively closer to the boundary of the system, which confirms the starting-line hypothesis we proposed.
VI Summary
For one-dimensional chain models where both hoppings and onsite potential are quasiperiodically modulated, we have analyzed their delocalization-localization transitions under equilibrium and non-equilibrium conditions, thermodynamic applications, and dynamic evolution behaviors. We numerically calculated and obtained the delocalization-localization phase diagram of the model, and verified it through MIPR and fractal dimensions. The results show that the delocalized phase with extended feature and and the localized phase can be separated by an ellipse curve. By using the model as the working medium of the quantum Otto cycle, we analyzed the underlying working modes of the cycle under multiple sets of typical parameters. The results show that when the system is in the delocalized phase with extended feature, it is conducive to the realization of the quantum heat engine, while being in the localized phase is conducive to the realization of the quantum heater. Furthermore, regardless of the phase region where the system is located, the working parameter space of the accelerator mode is always the largest, while that of the refrigerator mode never exists. After calculating the delocalization-localization phase diagram of the non-equilibrium steady state, we discussed the influence of quasiperiodic bond dissipation on the initial state evolution of the model and found that under moderate $V$ and $t_{1}$, the steady state of the system can largely maintain the initial delocalization property of the system. By adjusting $\gamma$ and $t_{1}$ or by adjusting the parameter $V$, the properties of the steady state can be regulated, gradually changing it from a delocalized state to a localized state. When the value of $V$ is much smaller, even if the initial state is localized, the steady state can also be delocalized. We analyzed the non-equilibrium dynamic evolution behavior and the existence of the quantum Mpemba effect using the trace distance. The results told that the quantum states with localized and critical properties as initial states are necessary conditions for the occurrence of the quantum Mpemba effect. These results can be explained by the starting-line hypothesis we have proposed. As for whether this hypothesis still holds true for other open quasi-periodic systems, this is an open question that remains to be tested.

### v3

@@ III Dissipation driven delocalization-localization transition
Recent research efforts have yielded conflicting perspectives regarding the steady-state delocalization-localization transition under dissipative modulation Longhi (2024); Liu et al. (2024); Peng et al. (2024) . By employing the dissipation, one can precisely control the delocalization and localization properties of the system under non-equilibrium steady states. Therefore, it motivates us to study the delocalization and localization properties of the steady state after introducing the dissipation. Besides, some studies suggest this transition is independent of the system’s initial properties Liu et al. (2024). Alternative viewpoints argue that the system’s initial attributes can influence the transition. Different initial property may either shorten or extend the threshold required for the delocalization-localization shift to occur. We aim to study whether the dependence of system’s initial property will also occur in this system with diagonal and off-diagonal quasiperiodic modulations and to study how different initial properties affect the delocalization-localization transition of the steady state as well.
@@
Upon introducing this bond dissipation Lindblad operator $L_{n}=c^{{\dagger}}_{n}c_{n+\ell}$, generated following the ones in Ref.Yusipov et al. (2017); Vershinina, O. S. et al. (2017); Jiang et al. (2024); Yang et al. (2025); Liu et al. (2024) , then the time-evolution of the density matrix $\rho$ is dictated by the following Lindblad master equation
@@
Here, the dissipator $\mathcal{D}\left[L_{n}\right]$ is defined as $\mathcal{D}\left[L_{n}\right]=L_{n}\rho L^{{\dagger}}_{n}-\{L^{{\dagger}}_{n}L_{n},\rho\}/2$, and $\gamma_{n}=\gamma\cos({2\pi\alpha n})$ represents the tunable quasiperiodic bond dissipation strength with the incommensurate parameter $\alpha=\frac{\sqrt{5}-1}{2}$. The rationale for our selection of the quasiperiodic form for the key dissipation operator stems from the consideration of the following aspect: the initial key dissipation operator is made up of four terms, with two being dephasing terms Longhi (2024) and the remaining two being hopping terms. As indicated in Ref. Liu et al. (2024), the dephasing effect fails to induce the delocalization-localization transition of steady states. It has been theoretically and experimentally confirmed that quasiperiodic modulation is capable of causing localization Roati et al. (2008). Moreover, this bond dissipation operator can be experimentally realized by introducing the auxiliary lattice Diehl et al. (2008, 2010, 2011). Owing to these factors, we utilized such a dissipative operator. This evolution equation can also be equivalently expressed as
@@
With $L=100$ and $\ell=1$, the delocalization phase diagrams for $V=0$, $V=0.5t$, $V=t$, and $V=2t$ are plotted in Figs. 3(a)- 3(d), respectively. The phase diagram is obtained by analyzing the position of the centroid in a steady state. We define a steady state where the centroid position (its definition is given in Eq.(9)) is no more than five times the lattice constant from the system boundary as a localized state; otherwise, it is a delocalized state. Our choice of criterion is guided by the experiment reported in Ref. Roati et al. (2008), which realized a one-dimensional Aubry–André (AA) model in a cold-atom optical lattice and observed the Anderson-localization transition; in the localized regime, the wave packet of an initially localized state broadened by approximately 10 lattice constants. Therefore, it is feasible to distinguish the delocalization-localization transition of the steady state through the position of the center of mass.
@@
As can be seen from Figs. 3(a)-3(c), when $t_{1}$ is less than $0.5t$, there is a large parameter range that makes all the steady states delocalized. This indicates that under these parameters, the steady state can retain the initial delocalized property of the system. As $\gamma$ or $t_{1}$ increases, the steady state gradually changes from a delocalized state to a localized state, meaning that the properties of the steady state can be regulated by tuning $\gamma$ and $t_{1}$. Meanwhile, we observe that as $V$ gradually increases, the range of the delocalized region gradually shrinks. It is indicated that adjusting the parameter $V$ can also regulate the localization property of the steady state. For instance, when V is at a relatively small value, it can be seen from 3(a) and 3(b) that when $t_{1}>t$, there are still delocalized states within it, resulting in the reentrance delocalization phenomenon. Moreover, the value of $V$ is smaller, making the delocalization region larger. When $V$ is relatively large, it can be seen from Figs. 3(c) and 3(d) that the delocalized region disappears. In particular, we find that when $V=2t$, even if the initial system is in the localized phase, the steady state is still delocalized (as can be seen from Fig. 3(d), the delocalized state exists within a small range of $t_{1}$ and $\gamma$).
IV quantum Mpemba effect and starting-line hypothesis
@@
Regarding the delocalization-localization transition in open quasiperiodic system, the focus on this topic mainly lies in the localization properties of steady states. For an anomalous dynamic process: the relaxation rate of a system whose initial state is further from the equilibrium state is actually faster than that of a system whose initial state is closer to the equilibrium state, that is, the quantum Mpemba effect, remains to be studied and discussed.
@@
Figure 4: (Color Online) Trace distance $d_{\rm Tr}$ versus evolution time $\tau$. (a) $V=0$, $t_{1}=1.2t$, and $\gamma=0.5t$. (b) $V=0$, $t_{1}=1.7t$, and $\gamma=0.8t$. (c) $V=0.2t$, $t_{1}=1.2t$, and $\gamma=0.2t$. (d) $V=0.2t$, $t_{1}=1.5t$, and $\gamma=0.2t$. (e) $V=0.2t$, $t_{1}=0.8t$, $\gamma=0.2t$. (f) $V=0.5t$, $t_{1}=0.5t$, $\gamma=0.2t$. The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$. Other parameter is $L=100$.
@@
We employ the density matrix $\rho(\tau)$ that describes the evolution of the system over time and the steady state density matrix $\rho_{ss}$ to calculate the standard distance between the matrices, this is, the trace distance, to characterize the quantum Mpemba effect. The trace distance $d_{\rm Tr}$ is defined as
@@
where $T=\rho(\tau)-\rho_{ss}$. Under Lindbladian dynamics, this quantity is monotonic in time and has been employed experimentally to probe both the strong quantum Mpemba effect and its inverse counterpart Aharony Shapira et al. (2024); Zhang et al. (2025). A nonzero trace distance $d_{\rm{Tr}}>0$ indicates that the system has not yet relaxed to the steady state, whereas $d_{\rm{Tr}}=0$ certifies convergence to the steady state.
Before examining the quantum Mpemba effect, we introduce the starting-line hypothesis. Consider a set of initial states characterized by their energies and center-of-mass (COM) positions. Let $E_{\mathrm{ss}}$ and $x_{\mathrm{com}}^{\mathrm{ss}}$ denote the energy and COM of the steady state, and define the COM distance as
$\Delta x_{\mathrm{com}}\equiv\left|x_{\mathrm{com}}(0)-x_{\mathrm{com}}^{\mathrm{ss}}\right|,$ (7)
[…]
where $n$s represent as the marks of the site index. The hypothesis posits: if the state whose energy is close to $E_{\mathrm{ss}}$ has a larger $\Delta x_{\mathrm{com}}$ than another state whose energy lies farther from $E_{\mathrm{ss}}$, then the latter can relax faster, manifesting the quantum Mpemba effect. This yields a simple criterion for which initial states can trigger the effect. For extended states, $x_{\mathrm{com}}(0)$ is pinned to the system center, so candidate extended initial states share almost the same ’starting line’ and $\Delta x_{\mathrm{com}}$ do not exhibit the effect. By contrast, localized and critical states generally have $x_{\mathrm{com}}(0)$ displaced from the center; such ’starting-line discrepancies’ enable a state that is farther in energy (but closer in COM) to relax faster than a state that is closer in energy (but farther in COM), thereby producing the quantum Mpemba effect.
@@
We take three different types of system eigenstates as the initial states and calculated the evolution of trace distance over time, as shown in Figures 4(a)-4(f). The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$. Specifically, the initial state energy corresponding to the red curve is closer to the steady-state energy than that of other initial states. Figures 4(a) and 4(b) show the results of the delocalized states with critical properties as the initial state. The COM positions corresponding to the initial states (rounded off, the same below) are respectively: $n=\{60,49,39\}$ and $n=\{60,49,41\}$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=\{57,47,36\}$ and $\Delta x_{\mathrm{com}}=\{56,46,38\}$. It can be seen from the data that the COM distance $\Delta x_{\mathrm{com}}$ between the initial state with energy close to the steady state and the steady state is actually greater than that when the energy is far from the steady state. The time required for $d_{\rm Tr}$ to evolve to zero is longer, and thus the so-called quantum Mpemba effect emerges.
Figures 4(c) and 4(d) show the results of the localized states as the initial states, where the corresponding COM positions of the initial states are respectively: $n=\{63,61,30,83,49\}$ and $n=\{66,63,15,56,71\}$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=\{60,58,27,80,46\}$ and $\Delta x_{\mathrm{com}}=\{61,58,10,51,66\}$. We can observe that when energy moves away from the initial state of steady-state energy, it evolves to the steady state more quickly compared to the quantum state where energy is close to the initial state energy. Therefore, the so-called quantum Mpemba effect also emerges. We can still observe that for the initial state where energy is far from the steady-state energy, it evolves to the steady state more quickly than the quantum state where energy is close to the steady-state energy. Therefore, the so-called quantum Mpemba effect emerges. This phenomenon can be explained by the proposed hypothesis as well, because the COM of the initial state with energy close to the steady state is further away from the steady state compared to other initial states, and thus it takes longer time to evolve to the steady state.
Figures 4(e) and 4(f) present delocalized states with extended properties as the results of initial states, where the corresponding COM positions of the initial states are respectively: all $n=50$ and all $n=51$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=42$ and $\Delta x_{\mathrm{com}}=21$. We can see that regardless of whether the energy is close to or far from the initial state of the steady-state energy, the time it takes for $d_{\rm Tr}$ to evolve to zero is almost indistinguishable, and thus the so-called quantum Mpemba effect does not occur. This phenomenon can be explained by using our hypothesis. Because the initial COM positions and COM distances of different initial states are almost the same, the time required for evolution to a steady state is also almost the same. Based on these results, we can conclude that for the initial states with extended properties, the dynamical evolution behavior does not exhibit the quantum Mpemba effect, while for the initial states with critical and localized properties, the dynamical evolution behavior will exhibit the quantum Mpemba effect. From the perspective of the COM position of the initial state, when the Mpemba effect occurs, the wave packet of the initial state is relatively closer to the boundary of the system, which confirms the starting-line hypothesis we proposed.
V Summary
We have investigated a one-dimensional tight-binding model with simultaneous quasiperiodic modulations of both hopping and onsite potential, focusing on localization–delocalization transitions under equilibrium and nonequilibrium conditions, and quantum relaxation dynamics. Using finite-size numerics, we have mapped the equilibrium phase diagram and verified it via the mean inverse participation ratio and fractal dimensions, finding that the delocalized (extended) and localized phases are separated by an approximately elliptical phase boundary. By computing the nonequilibrium steady-state phase diagram under quasiperiodic bond dissipation and analyzing initial-state dependence, we have known: for moderate $V$ and $t_{1}$, the steady state largely inherits the initial state’s delocalized character; tuning $\gamma$, $t_{1}$ or $V$ continuously drives the steady state from delocalized to localized, and for sufficiently small $V$ the steady state can become delocalized even from localized initial conditions. Finally, using the trace distance, we have demonstrated that there exists quantum Mpemba effect in which states initialized in localized or critical regimes can relax faster than states prepared closer to the steady state. These observations are rationalized by a starting-line hypothesis, whose applicability to broader classes of open quasiperiodic systems remains an open question.